# EvalLM: Interactive Evaluation of Large Language Model Prompts on User-Defined Criteria

### v1

@@ 5.1.2. Sampling Inputs and Generating Outputs
To test their prompts, the user can upload their own input dataset and then sample inputs from this dataset (Fig. 2C). Users are provided with two ways for sampling inputs: (1) manual, which opens a panel where the user can browse through their dataset and choose samples, and (2) diverse, which automatically samples distinct data points (details in §5.3). When the user samples inputs, these are shown as rows in the data panel, where the first column shows the content of the input sample (Fig. 5A). Then, the user can click on Run Prompts to generate outputs for each of these sampled inputs with each of their prompts. The outputs from each prompt are shown side-by-side in the data row (Fig. 5B).
@@ 5.1.3. Defining Criteria
The user defines new criteria by writing its name and providing a description for the criteria, which can explain what the LLM should assess when evaluating that criteria or describe the characteristics that an output must possess to satisfy that criteria (Fig. 4).
@@
This dictionary can serve as a starting point by providing initial criteria descriptions that the user can adapt to their context, or as inspiration by helping users consider other aspects to evaluate.
As Emily knows that explanations to children should only contain familiar words and information, she creates a criteria called “Familiarity” to assess whether generated examples use language and situations that a child can understand. Then, to decide on what else to evaluate, Emily browses through the dictionary and finds the “Faithfulness” criteria, which checks that summaries are devoid of factual errors (Krishna et al., 2023). As it is also important to verify that examples are faithful to the scientific information, Emily adds and edits the criteria to fit her context.
@@ 5.1.4. Revising Criteria: Refine, Merge, and Split
To improve the quality of their evaluations, users may need to iteratively revise their criteria.
To help users identify potential improvements in their criteria (DG4), EvalLM provides the Criteria Review Tool
that checks the criteria on their (1) clarity and relevance, (2) redundancy, and (3) granularity.
The tool automatically reviews the user’s criteria and, if it identifies criteria that could be improved, the system displays badges next to the criteria name (Fig. 4C) that the user can click to see the recommended improvements (Fig. 4D).
Similar to how external judges assess psychometric scales, the tool identifies criteria that are unclear or irrelevant, and suggests how to refine
the criteria to increase their clarity and relevance with the user’s task.
Inspired by item reduction analysis, the tool also identifies criteria that may be redundant and suggests how these could be merged
into one single criteria.
Finally, as prior work has demonstrated that human evaluators are more accurate when evaluating fine-grained criteria (Gehrmann et al., 2023; Krishna et al., 2023), the review tool identifies criteria that may be coarse-grained and suggests how to split
them into multiple fine-grained criteria.
@@
As she was reading through the input samples, Emily notices that the review tool has suggested that she splits the “Familiarity” criteria into “Language Simplicity” and “Relatability”. She agrees that, while these suggested criteria are related, they assess different aspects of the outputs. She decides to add these two suggestions and remove her previous “Familiarity” criteria.
@@ 5.1.5. Evaluating Outputs
After generating outputs and defining their criteria, the user can click on Auto-Evaluate to automatically evaluate the output pairs on each of the criteria (DG1). For each criterion, the evaluation assigns each output in a pair with a score out of 10 and, based on these scores, the interface presents which output was better at satisfying that criterion or if there was a tie (Fig. 5C). If the user wants to understand the evaluation results, they can click on the criteria name to view the explanation for that evaluation and the scores assigned to each output (Fig. 5D). To help the user verify the explanations without dedicating substantial effort into fully reading the outputs (DG2), the interface highlights fragments from the outputs (Fig. 5E) that the evaluation assistant focused on more when evaluating that criterion.
To provide a bigger picture on the performance of the prompts, the interface also displays statistics of the evaluation results (Fig. 2E). For each criterion, the interface presents the proportion of samples where each prompt “won” and the proportion of “ties”.
@@
After generating and evaluating outputs on three samples, Emily checks the overview of results and sees that her improved prompt won the most in all three criteria. However, she notices that there was one sample where the original prompt won in the “Language Simplicity” criteria. To check why this happened, she looks for the sample and opens the explanation for that evaluation, which describes that the output from her prompt used more complex terms. To improve on this, Emily adds an additional requirement to her prompt that instructs the LLM to avoid using complex terms and to always simplify them first.
@@
As the evaluations are conducted by a non-deterministic LLM, the evaluation results may differ in every run. To increase their assurance on the evaluation results, the user can increase the number of evaluation trials that are performed. If the number of evaluation trials is set to three, the interface evaluates each output pair three times and decides on the “winner” for each criteria based on which output won the most number of trials (i.e., majority vote). The user can check the evaluations for each of these trials by opening the explanation box and using the carousel at the bottom (Fig. 5F).
@@
5.1.6. Additional Features: History and Validation
To help the user keep track of their iterations, EvalLM automatically records all evaluations (DG2).
The user can click on the “History” button
(Fig. 2E) to view a visualization of their evaluation history, and look back at prompts and criteria used in those evaluations (Fig. 7).
Additionally, the system allows users to store generated samples into a validation set where they can annotate their own ground-truth evaluations.
With a populated validation set, the user can click on Validate Criteria (Fig. 2F) to evaluated these samples and check how accurately the automatic evaluator predicts the user’s own ground-truth evaluations.
5.1.7. Experimenting on Larger Samples
After designers have developed their prompts and criteria, they may wish to verify their prompt’s performance by testing on significantly larger samples. EvalLM provides an Experiment screen (full figure in Appendix A), where the user can automatically generate and evaluate larger samples and check the reliability of the evaluations. In this screen, designers can increase the number of evaluation trials and select another LLM as an alternative evaluator (e.g., ChatGPT or PaLM 2 (Anil et al., 2023)). When the designer runs an experiment, the system automatically samples diverse inputs, generates outputs with these inputs and the chosen prompts, and then evaluates the outputs on the chosen criteria for the configured number of trials. If an alternative evaluator was selected, this LLM is used to evaluate the same outputs on the same criteria for the same number of trials. The Experiment screen shows two additional statistics (Fig. 7): test-retest reliability or the consistency of evaluations between trials, and inter-rater reliability or the consistency between the evaluations by the system’s LLM (i.e., GPT-4) and the alternative evaluator. These statistics are shown as stacked bar charts that present the proportion of samples where evaluations were consistent or not, and the user can click on a bar to only display these unreliable cases (DG5).
After multiple iterations, Emily moves to the Experiment screen to choose between two promising prompts. After running an experiment with 40 samples and ChatGPT as the alternative evaluator, Emily sees that her first prompt excels at “Language Simplicity” while the second outperforms in “Scientific Accuracy”, representing a trade-off between the criteria. Additionally, she notices that the evaluators disagree often in “Language Simplicity” so she clicks on that bar to browse through these cases. By browsing through, Emily notices that the evaluation assistant may be unreliable as it also assessed sentence complexity for this criteria. As she wants it to focus only on vocabulary, she changes the criteria to “Simple Vocabulary”.
@@ 5.2.1. Automatic Evaluation
The main goal of EvalLM is to evaluate prompt outputs by comparing them on their ability to satisfy a set of user-defined criteria. To this end, we designed our evaluation prompt by adapting the prompt design from two state-of-the-art approaches for LLM-based evaluation: LLM-as-a-judge (Zheng et al., 2023), which compares model responses on their overall quality, and FLASK (Ye et al., 2023), which rates a single model response on how it performs on multiple “skills” or criteria that are described in natural language. Our prompt takes as input a task instruction, an input sample, a pair of outputs, and a list of criteria descriptions. Then, our prompt instructs an LLM to evaluate the output pair on each criterion by (1) explaining how the outputs satisfy the criterion, extracting evidence fragments from each output, and then providing each output with a score out of 10. To design this prompt, we considered alternative approaches for each of these components.
@@
Instruction and Input: We only include the overall instruction that is shared between the generation prompts to provide the evaluator LLM with context about the task while keeping its evaluations prompt-agnostic to limit potential bias due to the content or wording of each generation prompt.
@@
Output Ordering: Prior work has found that LLM-based evaluations have positional bias (Wang et al., 2023a) where they frequently favor the first output. While this work recommended evaluating each candidate in each position and aggregating these evaluations, as this would introduce additional cost and delays, our system instead alternates the positions of outputs every time it evaluates.
@@
Criteria Descriptions: Similar to Ye et al. (Ye et al., 2023), the criteria are added to the prompt as a list of lines of the form: “name of criterion: description of the criterion.” However, in their work, they also designed scoring rubrics for each criteria that describe the meaning of each numeric score. In contrast, we opt for only including the criteria descriptions as designing rubrics may require excessive effort from users and, according to our preliminary tests, automatically generated rubrics could negatively affect evaluation results.
@@
Explanation: The prompt instructs the LLM to provide an explanation where it compares and contrasts between the outputs as this explanation is later shown to users, but also because this could elicit reasoning from the LLM and increase performance (Wei et al., 2022). While we considered an initial design where the LLM was instructed to describe the performance of each output separately as this could lead to more in-depth examinations, initial tests and pilot studies revealed that this produced repetitive and less useful explanations.
@@
Evidence: To help users associate the evaluation explanations with the outputs, the prompt instructs the LLM to extract fragments from the outputs that are relevant to its explanation and evaluation. While we considered a design where the LLM first extracted evidence and then cited these in its explanation, we observed that the resulting explanations would simply list these fragments without providing in-depth explanations of how the outputs satisfied the criteria.
@@ 5.2.2. Criteria Review
To automatically review the user’s criteria and provide suggested revisions, we designed three prompts for each type of supported review: refining, merging, and splitting. Instead of designing one prompt that conducts all of these reviews, we designed separate prompts for each review as certain criteria may require multiple revisions and we wanted to allow users to flexibly choose between these. These three prompts follow the same general design: given a task instruction and a set of criteria, the LLM is instructed to review each criterion and identify any faulty ones, explain how these can be revised, and then return new criteria that result from these revisions. The three prompts differ in terms of what determines a criteria to be faulty and how they should be revised:
@@
Splitting: Identifies criteria that are excessively broad and consider multiple unrelated aspects, and divides these into more specific and mutually exclusive criteria.
@@ 5.3. Implementation Details
We implemented the front-end of EvalLM using TypeScript, ReactJS, and CSS. The back-end was implemented as a Flask server and the OpenAI API222https://platform.openai.com/ for all LLM components. In terms of the LLM settings, we set the temperature to 0.3 for all components. The automatic evaluation and criteria review tool used the gpt-4-0613 model and, as LLMs are prone to self-enhancement bias where it rates its own outputs highly (Zheng et al., 2023; Li et al., 2023), we use gpt-3.5-turbo-0613 when generating outputs. Finally, to support diverse sampling of inputs, we automatically cluster samples in the uploaded datasets by embedding data points using the OpenAI API with the text- embedding-ada-002 model, and then clustering these embeddings using the KMeans algorithm in the scikit-learn333https://scikit-learn.org/ library. When the user chooses to sample diversely, the system chooses data points from distinct clusters.
@@ 6. Technical Evaluation
We conduct a small-scale technical evaluation of our LLM-based evaluation approach to understand how performance is affected with more task-specific criteria, and to gain a more in-depth understanding of the LLM’s explanations for its evaluations.
@@ 6.1.1. Dataset
We assess LLM evaluations by comparing them to human evaluations in the MT-Bench dataset (Zheng et al., 2023). This dataset presents 80 user requests of diverse categories (e.g., writing, roleplay, math, coding) and the responses from various LLMs to each request. These responses are paired and, for each pair, the dataset provides votes from one to three human annotators on what response was better or if there was a tie. For our evaluation, we selected 19 requests in the writing and role-playing category as they involve the most subjectivity. As our evaluations focus on prompts rather than requests, we decompose these requests into a prompt-input format. We include examples of the original requests and their prompt-input versions in Appendix C, and the full data in the Supplementary Materials.
@@ 6.1.2. Conditions
Overall-Quality: We adopted the prompt from LLM-as-a-judge (Zheng et al., 2023) that compares outputs on overall quality.
General-Criteria: We used our evaluation prompt with general criteria from FLASK (Ye et al., 2023). In this work, they instruct an LLM to select three criteria (out of their set of 12) that are the most relevant to a given request. We followed their approach to select three criteria for each of the 19 tasks.
Specific-Criteria: We used our evaluation prompt with criteria that were automatically adapted for each task. For each task, we start with the same criteria from the General- Criteria condition, but automatically split and refine them using our criteria review prompt (i.e., all automatic suggestions are taken).
@@ 6.1.4. Results
Overall, Specific-Criteria had the highest agreement and correlation with human annotators and General-Criteria had the lowest (Table 1). By qualitatively reviewing cases, we observed that the Specific-Criteria condition could lead to more balanced evaluations compared to the other conditions. For example, when assessing the quality of generated travel blogs, Overall-Quality only assessed how many diverse attractions were covered, while Specific-Criteria also assessed the level of detail in which attractions were covered. Additionally, compared to General-Criteria, we observed how Specific-Criteria assessed various specific requirements that were posed by the prompt, enabling it to better capture how well the outputs followed the prompts. We also found cases where Specific-Criteria was less successful as each criterion was given equal weighting but certain criteria should hold precedence—e.g., not providing an incorrect medical diagnosis should be more important than the readability of an output. These preliminary findings suggest that instructing LLMs to evaluate more fine-grained and specific criteria can increase their ability to assess the alignment of outputs with instructions.

### v2

@@ 5.1.2. Sampling Inputs and Generating Outputs
To test their prompts, the user can upload their own input dataset and then sample inputs from this dataset (Fig. 2C). Users are provided with two ways for sampling inputs: (1) manual, which opens a panel where the user can browse through and choose samples, and (2) diverse, which automatically samples distinct data points (details in §5.3). When the user samples inputs, each one is shown as a row in the data panel (Fig. 5A). Then, the user can click on Run Prompts to generate outputs for each sampled input with each of their prompts. The outputs from each prompt are shown side-by-side in the data row (Fig. 5B).
@@ 5.1.3. Defining Criteria
The user defines a new criterion by providing a name and a description, which explains what the criteria assesses or the characteristics that an output must possess to satisfy that criteria (Fig. 4).
@@
This dictionary can serve as a starting point by providing initial descriptions that the user can adapt to their context, or as inspiration by helping users consider other aspects to evaluate.
Emily first creates a criterion, “Familiarity”, to check that generated examples only use language and situations that a child can understand. To decide on what else to evaluate, Emily browses through the dictionary and finds the “Faithfulness” criteria, which checks that summaries are devoid of factual errors (Krishna et al., 2023). Emily adds and edits this criterion to assess that generated examples are faithful to the given scientific information.
@@ 5.1.4. Revising Criteria: Refine, Merge, and Split
To help users identify potential improvements in their criteria and improve the quality of evaluations (DG4), EvalLM provides the Criteria Review Tool
that checks criteria for (1) clarity and relevance, (2) redundancy, and (3) granularity.
This LLM-powered tool automatically identifies criteria that could be improved (Fig. 4C) and recommends improvements (Fig. 4D).
Similar to how external judges assess psychometric scales (Robinson, 2018; Boateng et al., 2018), the tool identifies criteria that can be clearer or more relevant to the user’s task and suggests how to refine them
Inspired by item reduction analysis (Boateng et al., 2018), the tool also identifies criteria that may be redundant and suggests how these could be merged
into a single criteria.
Finally, as human evaluations are more accurate with fine-grained criteria (Gehrmann et al., 2023; Krishna et al., 2023), the review tool identifies coarse-grained criteria and suggests how to split
them into multiple criteria.
@@
As she was reading through outputs, Emily notices that the review tool has suggested splitting the “Familiarity” criteria into “Language Simplicity” and “Relatability”. As she agrees that these assess different aspects, she adds these two suggestions and removes her previous “Familiarity” criteria.
@@ 5.1.5. Evaluating Outputs
After generating outputs and defining criteria, the user can click on Auto-Evaluate to automatically evaluate the output pairs (DG1). For each criterion, the evaluation assigns each output with a score out of 10 to decide which output was better at satisfying that criterion or if there was a tie (Fig. 5C). If the user wants to understand the evaluations, they can click on a criteria name to view the explanation for that evaluation (Fig. 5D). To help the user verify the explanations without fully reading the outputs (DG2), the interface highlights fragments from the outputs (Fig. 5E) that were focused on more when evaluating the criterion. To provide a bigger picture on prompt performance, the interface also displays the proportion of samples where each prompt “won” and the proportion of “ties”. (Fig. 2E).
@@
After running the evaluation, Emily checks the results overview and sees that her improved prompt won the most in all three criteria, but lost once in the “Language Simplicity” criteria. To check why, she opens the evaluation explanation, which mentions that the improved prompt’s output used more complex terms. To improve on this, Emily adds a requirement to her prompt to always simplify complex words first.
@@
As LLMs are non-deterministic, the evaluation results may differ in every run. To increase their assurance of the evaluation results, the user can increase the number of evaluation trials. The interface will then evaluate each output pair for the chosen number of trials and decide on the “winner” for each criterion based on which output won the most number of trials (i.e., majority vote). The user can check the evaluations for each trial by using the carousel at the bottom of the evaluation explanations (Fig. 5F).
5.1.6. Additional Features: History and Validation
To help the user keep track of their iterations, EvalLM automatically records all of the user’s prompts, criteria, and evaluations (DG2, Fig. 6), which can be viewed by clicking on the “History” button
(Fig. 2E).
Additionally, users can store generated samples into a validation set where they can annotate their own ground-truth evaluations.
With a populated validation set, the user can Validate Criteria (Fig. 2F) to assess how accurately the automatic evaluator predicts the user’s ground-truth evaluations.
5.1.7. Experimenting on Larger Samples
After designers have developed their prompts and criteria, they may wish to verify their prompt’s performance by testing on significantly larger samples. For this purpose, EvalLM provides the Experiment screen (full figure in Appendix A). In this screen, designers can set the number of evaluation trials and select an alternative evaluator LLM (e.g., ChatGPT or PaLM 2 (Anil et al., 2023)). When the user runs an experiment, the system automatically samples diverse inputs, generates outputs with these inputs and the chosen prompts, and then evaluates the outputs on the chosen criteria for the configured number of trials. If an alternative evaluator was selected, this LLM is used to evaluate the same outputs on the same criteria for the same number of trials. The Experiment screen shows two additional statistics (Fig. 7): test-retest reliability or the consistency of evaluations between trials, and inter-rater reliability or the consistency between the evaluations by the system’s LLM (i.e., GPT-4) and the alternative evaluator. These statistics are shown as stacked bar charts that present the proportion of consistent and inconsistent evaluations, and the user can click on a bar to only display those cases (DG5). In this screen, the user can also click on the stacked bar charts for the evaluation overview to see only the samples where one prompt performed better than the other (or they were tied) for a chosen criteria.
To decide between two promising prompts, Emily runs an experiment with 40 samples and ChatGPT as the alternative evaluator. The results show that her first prompt excels at “Language Simplicity” but loses in “Scientific Accuracy”, signaling at a possible trade-off. Emily also notices that the evaluators often disagree in “Language Simplicity” and, by clicking on that bar to browse through cases, she finds that GPT-4 also assesses sentence complexity for the criteria. As she only wants it to assess vocabulary, she changes it to “Simple Vocabulary”.
@@ 5.2.1. Automatic Evaluation
The main goal of EvalLM is to evaluate prompt outputs by comparing them on a set of user-defined criteria. To this end, we designed our evaluation prompt by adapting the prompts from two state-of-the-art approaches: LLM-as-a-judge (Zheng et al., 2023), which compares model responses on their overall quality, and FLASK (Ye et al., 2023), which rates the performance of a single response on multiple “skills” or criteria. Our prompt takes as input: a task instruction, an input sample, a pair of outputs, and a list of criteria descriptions. Then, our prompt instructs an LLM to evaluate the output pair on each criterion by (1) explaining how the outputs satisfy the criterion, (2) extracting evidence fragments from each output, and then (3) providing each output with a score out of 10. To design this prompt, we considered alternative approaches for each of these components.
@@
Instruction and Input: We only include the overall instruction that is shared between the generation prompts to provide the evaluator LLM with context about the task while keeping its evaluations prompt-agnostic—limiting potential bias due to the content or wording of each generation prompt.
@@
Output Ordering: Prior work found that LLM-based evaluations have positional bias, frequently favoring the first output (Wang et al., 2023a). While this work suggested evaluating each candidate in each position and aggregating results, this would introduce additional costs and delays. Instead, our system alternates the positions of outputs in every evaluation.
@@
Criteria Descriptions: Similar to Ye et al. (Ye et al., 2023), the criteria are added to our prompt as lines of the form: “name of criterion: description of the criterion.” While they also included scoring rubrics for each criterion to describe each numeric score, we only included the criteria descriptions. We considered that designing rubrics would require excessive effort from users and, based on preliminary tests, generated rubrics could negatively affect evaluations.
@@
Explanation: The prompt instructs the LLM to provide an explanation where it compares and contrasts between the outputs. This explanation is useful for users, but can also elicit reasoning from the LLM and increase performance (Wei et al., 2022). While we considered a design where the LLM explains the performance of each output separately to yield more in-depth examinations, pilot studies showed that these explanations were repetitive and less useful.
@@
Evidence: To help users associate the evaluation explanations with the outputs, we instruct the LLM to extract fragments from the outputs that are relevant to its evaluation. We also considered a design where the LLM first extracts evidence and then cites these in its explanation, but saw that these explanations would simply list the evidence without elaborating on them.
@@ 5.2.2. Criteria Review
To automatically review the user’s criteria and suggest revisions, we designed three prompts for each type of supported review: refining, merging, and splitting. Instead of designing one prompt that conducts all of these reviews, we designed separate prompts for each review as certain criteria may require multiple revisions and we wanted to allow users to flexibly choose between these. These three prompts follow the same general design: given a task instruction and set of criteria, the LLM is instructed to review each criterion and identify any faulty ones, explain how these can be revised, and then generate new criteria that result from these revisions. The three prompts differ in terms of what determines a criterion to be faulty and how they should be revised:
@@
Splitting: Identifies criteria that are excessively broad and consider multiple unrelated aspects, and divides these by generating new criteria for each of these aspects.
@@ 5.3. Implementation Details
We implemented the front-end of EvalLM using TypeScript, ReactJS, and CSS. The back-end was implemented as a Flask server and the OpenAI API222https://platform.openai.com/ for all LLM components. In terms of the LLM configurations, we set the temperature to 0.3 for all components. The automatic evaluation and criteria review tool used the gpt-4-0613 model and, as LLMs are prone to self-enhancement bias where it rates its own outputs highly (Zheng et al., 2023; Li et al., 2023), we use gpt-3.5-turbo-0613 when generating outputs. Finally, to support diverse sampling of inputs, we automatically cluster samples in the uploaded datasets by embedding data points using the OpenAI API with the text- embedding-ada-002 model and clustering these embeddings using the KMeans algorithm in the scikit-learn333https://scikit-learn.org/ library. To sample diversely, the system chooses data points from distinct clusters.
@@ 6. Technical Evaluation
We conduct a small-scale technical evaluation of our LLM-based evaluation approach to understand how performance is affected by more task-specific criteria, and to gain a more in-depth understanding of the LLM’s explanations for its evaluations.
@@ 6.1.1. Dataset
We assess LLM evaluations by comparing them to human evaluations in the MT-Bench dataset (Zheng et al., 2023). This dataset presents 80 user requests of diverse categories (e.g., writing, roleplay, math, coding) and responses from various LLMs to each request. These responses are paired and, for each pair, the dataset provides votes from one to three human annotators on what response was better or if there was a tie. For our evaluation, we selected 19 requests in the writing and role-playing categories as they involve the most subjectivity. As our LLM evaluations focus on prompts rather than requests, we decompose these requests into a prompt-input format. We include examples of the original requests and their prompt-input versions in Appendix C, and the full data in the Supplementary Materials.
@@ 6.1.2. Conditions
Overall-Quality: We adopted the prompt from LLM-as-a-judge (Zheng et al., 2023) that compares a pair of outputs to select the one with higher overall quality.
General-Criteria: We used our evaluation prompt with the more general and broad criteria from FLASK (Ye et al., 2023). In this work, they instruct an LLM to select three criteria, out of a set of 12, that are the most relevant to a given request. Then, this condition evaluated a pair of outputs to determine which one was better at satisfying each of the criteria.
Specific-Criteria: We used our evaluation prompt with criteria that were automatically refined and adapted for each request or prompt. For each request, we start with the same criteria that were selected by the LLM in the General- Criteria condition, but we automatically split and refine them using our criteria review technique (i.e., all automatic suggestions are taken). Then, a pair of outputs was evaluated by determining which better satisfied each of these more fine-grained and specific criteria.
@@ 6.1.4. Results
Overall, we observed that Specific-Criteria had the highest agreement and correlation with human annotators and General-Criteria had the lowest (Table 1). As a reference, for data points with at least two annotators, the Fleiss’ Kappa between two random human annotators was 0.496 (sampled 5 times and averaged), showing that Specific-Criteria agreed with human evaluations to a degree that was similar to human-human agreement. By qualitatively reviewing cases, we observed that the Specific-Criteria condition could produce more balanced evaluations. For example, when assessing generated travel blogs, Overall-Quality only assessed the breadth of attractions covered, while Specific-Criteria also assessed the depth of the attraction descriptions. Also, compared to General-Criteria, we saw that Specific-Criteria assessed specific requirements that were posed in prompts, enabling it to better capture how well outputs followed the prompts. We also found cases where Specific-Criteria was less successful as each criterion was given equal weighting, but certain criteria might need to hold precedence—e.g., not providing an incorrect medical diagnosis is more important than readability. These preliminary findings suggest that instructing LLMs to evaluate more fine-grained and specific criteria can increase their ability to assess the alignment of outputs with instructions. However, we also note that the agreement between LLM and human evaluations is still not perfect. While this could be attributed to the subjectivity involved, it also highlights the limitations of only relying on LLM evaluations.